OpenStudy (studygurl14):

Hello! Welcome to StudyGurl14's Study Session! Feel free to ask any questions and I'll try my best to walk you through it and give you an answer. :) Satisfaction guaranteed! (Maybe...lol) Medals, fanning, and/or testimonies are appreciated if you receive the answer you are looking for, but not in any way required. Thanks and let the questions begin! ~StudyGurl14

OpenStudy (anonymous):

ABCD is a rectangle with diagonals AC and BD intersect at E. If AE=x+2 and BD= 4x-16. Find the length of the diagonals of rectangle? Pls help

OpenStudy (rianiscool):

Which property justifies this statement? r(s + t) = rs + rt A. associative B. distributive C. symmetric D. transitive

OpenStudy (rianiscool):

lol I'm back

OpenStudy (anonymous):

solve each inequality. -6w < 84 w>14 w<-14 w>-14 w<14

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

@rian distributive, which states that a(b=c)= ab+ ac

OpenStudy (rianiscool):

:( don't answer let me try

OpenStudy (anonymous):

@rianiscool sorry!

OpenStudy (rianiscool):

b

OpenStudy (anonymous):

@kitkatNat1316 @StudyGurl14 ABCD is a rectangle with diagonals AC and BD intersect at E. If AE=x+2 and BD= 4x-16. Find the length of the diagonals of rectangle? Pls help guys

OpenStudy (anonymous):

can anyone answer mine

OpenStudy (rianiscool):

is that right @kitkatNat1316

OpenStudy (causticsyndicalist):

*farts* Okay, whats going on here and who needs help from the omega wolf himself?

OpenStudy (anonymous):

@rianiscool yes, try these problems: 3(x+5) 5(7+y)

OpenStudy (anonymous):

@Caus lol

OpenStudy (studygurl14):

Sorry stupid computer problems...who hasn't been helped?

OpenStudy (anonymous):

@rianiscool you would need to use the distributive property, if that helps

OpenStudy (anonymous):

@CausticSyndicalist ABCD is a rectangle with diagonals AC and BD intersect at E. If AE=x+2 and BD= 4x-16. Find the length of the diagonals of rectangle? Pls help

OpenStudy (anonymous):

@StudyGurl14 Pls reply ABCD is a rectangle with diagonals AC and BD intersect at E. If AE=x+2 and BD= 4x-16. Find the length of the diagonals of rectangle?

OpenStudy (rianiscool):

so b was i right

OpenStudy (anonymous):

@rianiscool yes, good job!

OpenStudy (rianiscool):

thank you

OpenStudy (rianiscool):

Translate the word phrase into a math expression. nine more than five times a number A. 9 > 5n B. 9 – 5n C. 5n – 9 D. 5n + 9

OpenStudy (anonymous):

B

OpenStudy (rianiscool):

for who @Vegiti

OpenStudy (rianiscool):

is it a

OpenStudy (studygurl14):



OpenStudy (anonymous):

D, i mean

OpenStudy (anonymous):

@rianiscool well, think about it, more= what? plus or minus? Nine more than 5 times a number

OpenStudy (anonymous):

This post is confusing....

OpenStudy (anonymous):

@rianiscool , so 9+5n, because more is the same as adding on to something. Make sense?

OpenStudy (rianiscool):

no it is d

OpenStudy (rianiscool):

Answer is nine more than five times a number = 9+5n Therefore D is the answer

OpenStudy (anonymous):

Thats what i said....

OpenStudy (anonymous):

@rianiscool exactly, good job (:

OpenStudy (anonymous):

4x-16=x+2

OpenStudy (rianiscool):

yes thank you @Vegiti

OpenStudy (anonymous):

Anytime fella! :)

OpenStudy (studygurl14):

2(x + 2) = 4x - 16 2x + 4 = 4x - 16 -2x = -20 x = 10 4x + 16 4(10) + 16 40 + 16 The answer: 56 @PeepzStudyTime

OpenStudy (studygurl14):

Thank everyone for helping out! (since my computer is acting up...lol)

OpenStudy (anonymous):

@StudyGurl14 np :)

OpenStudy (causticsyndicalist):

I didn't do much but you're welcome

OpenStudy (anonymous):

@CausticSyndicalist haha

OpenStudy (anonymous):

@StudyGurl14 are u sure that this 56 is the answer?

OpenStudy (studygurl14):

yes, why

OpenStudy (rianiscool):

Which situation could be solved using this expression? 2k – 3 A. Tamara made three times as many free throws as Marla. Marla made 3 free throws. How many free throws did Tamara make? B. Tamara made 3 fewer than twice the number of free throws Marla made. How many free throws did Tamara make? C. Tamara made 2 free throws in the first half of the game. Marla made 3 fewer free throws than Tamara in the game. How many free throws did Marla make? D. Tamara made 2 free throws. Marla made 3 free throws. Jenny made some free throws. How many free throws did they make altogether?

OpenStudy (anonymous):

hey peep